OpenStudy (anonymous):

Write the standard form of the equation of the circle with radius r and center (h, k) r =√15; (h, k) = (0, -10)

OpenStudy (anonymous):

the standart form is (x-xc)²+(y-yc)²=r²

OpenStudy (anonymous):

so x²+(y+10)²=15

OpenStudy (anonymous):

\[x^2+(y+10)^2=15\]

OpenStudy (anonymous):

if you like my answer,click good answer ,ty